The Gourmand Cooking School runs short cooking courses at its small campus. Management has identified two cost drivers it uses in its budgeting and performance reports—the number of courses and the total number of students. For example, the school might run two courses in a month and have a total of 65 students enrolled in those two courses. Data concerning the company’s cost formulas appear below:

Fixed Cost per Month Cost per Course Cost per
Student
Instructor wages $ 2,980
Classroom supplies $ 290
Utilities $ 1,220 $ 55
Campus rent $ 4,800
Insurance $ 2,000
Administrative expenses $ 3,500 $ 42 $ 7

For example, administrative expenses should be $3,500 per month plus $42 per course plus $7 per student. The company’s sales should average $900 per student.

The company planned to run four courses with a total of 65 students; however, it actually ran four courses with a total of only 63 students. The actual operating results for September appear below:

Actual
Revenue $ 55,600
Instructor wages $ 11,200
Classroom supplies $ 18,700
Utilities $ 1,850
Campus rent $ 4,800
Insurance $ 2,140
Administrative expenses $ 3,549

Required:

1. Prepare the company’s planning budget for September.

2. Prepare the company’s flexible budget for September.

3. Calculate the revenue and spending variances for September.
Answer #1
1
Planning Budget
For the month ended September 30
Budgeted Courses 4
Budgeted Students 65
Revenue 58500 =65*900
Expenses:
Instructor wages 11920 =4*2980
Classroom supplies 18850 =65*290
Utilities 1440 =1220+(4*55)
Campus rent 4800
Insurance 2000
Administrative expenses 4123 =3500+(4*42)+(65*7)
Total expense 43133
Net operating income 15367
2
Flexible Budget
For the month ended September 30
Actual Courses 4
Actual Students 63
Revenue 56700 =63*900
Expenses:
Instructor wages 11920 =4*2980
Classroom supplies 18270 =63*290
Utilities 1440 =1220+(4*55)
Campus rent 4800
Insurance 2000
Administrative expenses 4109 =3500+(4*42)+(63*7)
Total expense 42539
Net operating income 14161
3
Actual results Revenue and spending variance Flexible budget
Courses 4 4
Students 63 63
Revenue 55600 1100 U 56700
Expenses:
Instructor wages 11200 720 F 11920
Classroom supplies 18700 430 U 18270
Utilities 1850 410 U 1440
Campus rent 4800 0 None 4800
Insurance 2140 140 U 2000
Administrative expenses 3549 560 F 4109
Total expense 42239 300 F 42539
Net operating income 13361 800 U 14161